You have $2,000 in an account which pays 5% annual interest.

How much additional dollars of interest would you earn over 6 years if you moved the money to an account earning 6%
Answer #1

Ans $ 156.85

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
At 5%
FV = PV (1 + r )n
FV = 2000*(1+5%)^6
FV = 2680.19
At 6%
FV = PV (1 + r )n
FV = 2000*(1+6%)^6
FV = 2837.04
Additional Interest = 2837.04 - 2680.19
156.85